OpenStudy (anonymous):

The National Fish and Wildlife Foundation has asked you to calculate when the population of the fish reaches 500. Show all of your work. This is my equation:P(x)=10*2^x Please help me...

OpenStudy (anonymous):

anybody?!

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

please could someone help me? @DanJS?

OpenStudy (danjs):

That function is population P at time x P(x) = 10* 2^x

OpenStudy (danjs):

What information is given in the problem?

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

The part that i get choked up on is number 5. 1.Create the data to fill in the population table. Be creative when selecting the P(x) values, but remember that this is an exponential growth model. Notice that you even need to come up with how many fish the fisherman introduces to the lake. Months (x) Population P(x) 0 10 1 20 2 40 3 80 2.Describe how to create the exponential function represented by your table. Use complete sentences and arrive at a final exponential function, P(x). 3.Using your exponential function, P(x), explain the meaning behind each of the numbers and variables in terms of the invasive fish species. 4.The current exponential function P(x) measures the fish population’s growth monthly. Using complete sentences explain how to find the rate of growth every week. 5.The National Fish and Wildlife Foundation has asked you to calculate when the population of the fish reaches 500. Show all of your work. 6.The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm.

OpenStudy (danjs):

So you did #!1-4 already

OpenStudy (danjs):

and got that p(x) = 10*2^x ?

OpenStudy (danjs):

#1 The general exponential model is y = A*b^x you found A and B from the table right?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

lol okay!!

OpenStudy (danjs):

yeah, the data points fit in that equation, it is right

OpenStudy (anonymous):

So should i write 500=10*2^x?

OpenStudy (danjs):

P(x) = 10 * 2^x #5) When (what x), is it when P(x)=500 fish

OpenStudy (danjs):

yeah

OpenStudy (anonymous):

so how do i solve it?

OpenStudy (danjs):

ok, i will write words, you type out the equations....

OpenStudy (danjs):

divide both sides by 10 first

OpenStudy (anonymous):

500=10*2^x /10 /10 50=2^x

OpenStudy (danjs):

take log base 2 of everything

OpenStudy (danjs):

\[\log _{2}2^x = x\]

OpenStudy (anonymous):

i get 0?

OpenStudy (danjs):

i dont think you can use log base 2 in a calculator, we have to use the change of base formula

OpenStudy (anonymous):

wow,slow me

OpenStudy (danjs):

x = \[x = \log_{2}50 = \frac{ \ln(50) }{ \ln(2) }\]

OpenStudy (danjs):

remember that thing?

OpenStudy (anonymous):

yea. so i would get 5.64?

OpenStudy (danjs):

no

OpenStudy (danjs):

5.64 months

OpenStudy (danjs):

: )

OpenStudy (anonymous):

Thats it right?

OpenStudy (danjs):

yeah, the last question is already answered it looks like, but instead of using log base e, (ln), use log base 10

OpenStudy (danjs):

log base 10 is the 'LOG' button on the calculator

OpenStudy (danjs):

\[\frac{ \log_{10}50 }{ \log_{10} 2 }\]

OpenStudy (anonymous):

lol, i know that!!!

OpenStudy (danjs):

oh, beg my pard. lol

OpenStudy (anonymous):

i didnt wirte that^. computer is acting up.

OpenStudy (danjs):

Need to transform P(x) = 10 * 2^x to a log equation

OpenStudy (anonymous):

would it be like log2^x=10?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ p(x) }{ 10 } = 2^x\]\[\log_{2}[\frac{ p(x) }{ 10 }] = x\]

OpenStudy (danjs):

Then they want you to change that to base 10 log

OpenStudy (anonymous):

so you would change the 2 and 10 around?

OpenStudy (danjs):

\[x = \frac{ \log _{10}(p(x)/10) }{ \log _{10}2 }\]

OpenStudy (danjs):

the change of base formula thing

OpenStudy (anonymous):

yea thats what i mean.

OpenStudy (danjs):

that is the final answer i think

OpenStudy (danjs):

If you let p(x) be any number of fish, then you can use the computer, which can only do base 10 logs, to find how many months that will be

OpenStudy (anonymous):

okay!!!! I really appreciate your help DanJS!!! :) your the only one that helped me. really grateful for that :)

OpenStudy (anonymous):

you have a medal coming your way,lol

OpenStudy (danjs):

That is what i am here for.. you are welcome

OpenStudy (danjs):

I dont understand how people get confused about logarithms. Not just you, most people. They are just exponents. lol

OpenStudy (danjs):

\[\log_{b} x = y ~~~~or~~~~~b^y = x\] log base b of x, means ' b to what power equals x'

OpenStudy (anonymous):

lol, true i get confuse on what formula to use. and remembering the order of sequence, yea thats me!!!

OpenStudy (danjs):

not just you, 99% of people. Just think to hard . lol.

OpenStudy (anonymous):

yea i do that a lot. when i get the answer. i say to myself, man i knew that!!

OpenStudy (danjs):

Right, but the work to get to the answer is what has you remember things, if someone just told you the answer, you arent going to remember jack s--t

OpenStudy (anonymous):

SOOOO true!!!!! lmbo!!!!

OpenStudy (danjs):

I see math sorta like a toolbox of skills. You dont have to remember anything in particular, as long as you know the general landscape. So if you start somewhere, you can work towards something else using what you do remember.

OpenStudy (anonymous):

yea. i like your sense of thinking man.

OpenStudy (danjs):

: ) feel free to tag me in more if ya get stuck on anything

OpenStudy (anonymous):

will do!! thanks again DanJS!!! :)